# Modeling Elastic-Body Dynamics of Fish Swimming Using a Variational Framework

### v1

@@ Abstract
Fish-inspired aquatic robots are gaining increasing attention in research communities due to their high swimming speeds and efficient propulsion enabled by flexible bodies that generate undulatory motions. To support the design optimizations and control of such systems, accurate, interpretable, and computationally tractable modeling of the underlying swimming dynamics is indispensable. In this letter, we present a full-body dynamics model for fish swimming, rigorously derived from Hamilton’s principle. The model captures the continuously distributed elasticity of a deformable fish body undergoing large deformations and incorporates fluid–structure coupling effects, enabling self-propelled motion without prescribing kinematics. A preliminary parameter study explores the influence of actuation frequency and body stiffness on swimming speed and cost of transport (COT). Simulation results indicate that swimming speed and energy efficiency exhibit opposing trends with tail-beat frequency and that both body stiffness and body length have distinct optimal values. These findings provide insights into biological swimming mechanisms and inform the design of high-performance soft robotic swimmers.
@@ I Introduction
Biomimetic swimming robots have drawn increasing interest in research communities in recent years, due to their potential for energy-efficient and agile locomotion in complex aquatic environments [1, 2]. Among these biomimetic swimmers, fish-like robots are particularly promising, since they utilize flexible bodies and undulatory motion to achieve high speed, propulsive efficiency, and maneuverability [3, 4, 5]. Therefore, accurate modeling of the dynamics of fish swimming is essential, not only for understanding the underlying principles of aquatic locomotion, but also for insights into the design, optimizations, and control of bio-inspired robotic swimmers.
Plenty of theoretical and computational efforts have been made to study the dynamics of fish swimming. A common approach is to model the dynamics based on multi-segment methods [6, 7, 8, 9, 10, 11, 12]. Most of these methods model the fish as a chain of rigid bodies, and the elasticity of the fish’s body is lumped into the joints using torsional springs; though the work in [12] adopts an elastic rod framework, the actual implementation reduces the fish body to a chain of rigid segments with locally defined bending potentials limited to the connections of segments. An exception is the model presented in [10], where an arc is used to approximate the deformation of each segment and thus adds another degree of freedom for each segment.
@@
Another approach to studying fish swimming dynamics is to perform numerical simulations based on computational fluid dynamics (CFD) and fluid-structure interactions (FSI); the latter solves for both the solid and fluid fields. The FSI-based simulations presented in [13] successfully provides high-fidelity results of muscle-actuated fish swimming. A CFD-based model that accurately captures the vortex shedding during fish swimming is presented in [14]; however, the work prescribes the deformation pattern of the fish’s body, and thus does not provide a full-dynamics analysis. The work in [15] also presents an FSI-based model that provides a detailed analysis on the fluid dynamics of surrounding water and the effects of stiffness tuning on propulsion performances; however, the solid part of the model is reduced to a two-rigid-body linkage with a joint loaded by nonlinear spring and damper.
@@
Building on the preceding discussion, this letter presents a physics-grounded and fully interpretable model that captures the full-body dynamics of fish swimming. The model accounts for continuously distributed body elasticity with arbitrarily large deformations, incorporates both reactive and resistive hydrodynamic forces from the surrounding fluid, and enables self-propulsion without any imposed or prescribed kinematics. We further conduct a preliminary parameter study, investigating how swimming performance is influenced by actuation frequencies, body stiffness, and body length, thereby providing insights into both the fundamental dynamics of fish swimming and the design of efficient robotic swimmers.
@@ IV-C The Projected Weak Form
It is worth reiterating that all spatial integrals are evaluated using the Gauss-Legendre quadrature, and all derivatives (gradient, Hessian, and Jacobian computations) are evaluated using AD. The resulting ODEs can then be solved using numerical integrators that are accessible in most scientific computing languages. Since the mass matrix $\bm{M}$ is naturally ill-conditioned for this system given the strong coupling among the rigid head, the elastic body, and surrounding water, stiffness-aware methods (e.g., implicit methods) are needed for both efficiency and accuracy.
@@ V Results
To validate our model and to provide insights on fish swimming and potential future design of robotic fish, we perform convergence experiments for the Ritz series and a preliminary design analysis based on numerical simulation results.
We implement our numerical simulations in Julia, using the ForwardDiff.jl package for automatic differentiations and the DifferentialEquations.jl package for numerical integrations [26]. We mainly use two stiffness-aware numerical integrators from the package - RadauIIA5 (a stable and accurate solver that implements the implicit Runge-Kutta method) and TRBDF2 (an efficient solver that combines the trapezoidal rule with the second-order backward difference formula). Our Julia code runs on a Dell G15 laptop with a 13th Gen Intel(R) Core(TM) i7-13650HX 2.60 GHz CPU and 16 GB RAM.
In the design analysis, we sweep over three parameters respectively - the tail-beat frequency, the baseline Young’s modulus of elasticity of the fish’s body, and the length of the fish’s body, where the Young’s modulus distribution is modeled as
@@
where $E_{b}$ is the baseline Young’s modulus. The amplitude of the joint motor oscillations is fixed at $0.4\text{\,}\mathrm{rad}$ (i.e., the joint motor sweeps over a range of $45.86\text{\,}\mathrm{\SIUnitSymbolDegree}$). We focus on two criteria for evaluating the propulsion performances of the robotic fish - the forward speed and the cost of transport (COT), defined as
@@
where $W_{m}$ is the total work done by the joint motor, $m_{\text{total}}$ is the total mass of the robotic fish, $g$ is the gravitational acceleration, and $d$ is the total distance travelled. Since the fish undulates as it moves forward, we use the displacement magnitude in the forward direction (i.e., $X$-direction) as the effective travelled distance. The COT serves as an important metric for evaluating the efficiency of propulsion during fish swimming. We use a PD controller on the joint angle to control the frequency and amplitude of the joint motor.
@@ V-A Convergence of Ritz Series
The correct number of Ritz basis functions (i.e., the number of terms used in the Ritz series) at which the system shows convergent behavior need to be chosen. We set the baseline Young’s modulus to $0.35\text{\,}\mathrm{MPa}$, a typical Young’s modulus for bio-based soft materials [27]; we also set the tail-beat frequency to $2\text{\,}\mathrm{Hz}$ and joint motor amplitude to $0.4\text{\,}\mathrm{rad}$ (i.e., the joint motor sweeps over a range of $45.86\text{\,}\mathrm{\SIUnitSymbolDegree}$). We increase the number of Ritz basis functions from 1 to 6, and compare the simulated trajectories over 3 full tail-beats (i.e., $1.5\text{\,}\mathrm{s}$) from rest and the time consumed by the numerical simulations. The comparisons are shown in Fig. 3.
@@
V-B Parameter Sweep over Tail-Beat Frequency
We fix the body length to $0.3\text{\,}\mathrm{m}$ and the baseline Young’s modulus to $0.35\text{\,}\mathrm{MPa}$. We then perform a parameter sweep for the tail-beat frequency, ranging from $0.5\text{\,}\mathrm{Hz}$ to $3.5\text{\,}\mathrm{Hz}$ with an interval of $0.5\text{\,}\mathrm{Hz}$; the results are shown in Fig. 4.
@@ V-B Parameter Sweep over Tail-Beat Frequency
It can be observed from Fig. 4 that the steady-state forward speed peaks at a tail-beat frequency of $3\text{\,}\mathrm{Hz}$, while the COT is minimized at a tail-beat frequency of $1\text{\,}\mathrm{Hz}$.
V-C Parameter Sweep over Baseline Young’s Modulus
We fix the body length to $0.3\text{\,}\mathrm{m}$ and the tail-beat frequency to $2\text{\,}\mathrm{Hz}$. Then we perform a parameter sweep for the baseline Young’s modulus of the fish’s body, ranging from $0.2\text{\,}\mathrm{MPa}$ to $0.5\text{\,}\mathrm{MPa}$ with an interval of $0.05\text{\,}\mathrm{MPa}$; the results are shown in Fig. 5.
@@ V-C Parameter Sweep over Baseline Young’s Modulus
From Fig. 5, the forward speed is optimized at $0.25\text{\,}\mathrm{MPa}$ and COT is optimized at $0.3\text{\,}\mathrm{MPa}$.
V-D Parameter Sweep over Body Length
We fix the body Young’s modulus to $0.35\text{\,}\mathrm{MPa}$ and the tail-beat frequency to $2\text{\,}\mathrm{Hz}$. Then we perform a parameter sweep on the length of the fish’s body, ranging from $0.25\text{\,}\mathrm{m}$ to $0.45\text{\,}\mathrm{m}$ with an interval of $0.05\text{\,}\mathrm{m}$; the results are shown in Fig. 6.
Figure 6: Simulation results of the body length sweep. Time-lapse plot of the fish from $t=4s$ to $t=5s$ with a body length of (a) $0.25\text{\,}\mathrm{m}$ and (b) $0.45\text{\,}\mathrm{m}$ (blue-red-green-black sequence); (c) $X$-direction velocity plots and (d) plots of steady-state average speed and COT at different body lengths.
It can be observed from Fig. 6 that both the forward speed and the COT are optimized at a body length of $0.35\text{\,}\mathrm{m}$.
VI Discussion
From the numerical experiments, we observe that varying the tail-beat frequency has the most impact on both steady-state speed and the COT. However, the optimal frequencies for speed and COT do not occur simultaneously - for a tail-beat frequency of $\geq$1\text{\,}\mathrm{Hz}$\$, increasing the steady-state speed results in an increase in COT (which implies a decrease in propulsion efficiency), and vice versa. Therefore, trade-offs need to be made and priorities need to be taken.
It is also interesting to observe the non-convex COT distribution during the parameter sweep on the baseline Young’s modulus of the fish’s body. It could be a future direction of study to see if this shape holds for all neighboring tail-beat frequencies and body lengths. It is also important to note that, though varying the baseline Young’s modulus at the range of $0.2\text{\,}\mathrm{MPa}$ to $0.5\text{\,}\mathrm{MPa}$ has limited impact on both the steady-state speed and the COT, this does not imply that the impact of the Young’s modulus distribution throughout the fish’s body is trivial - in our numerical experiments, we assume that the Young’s modulus distribution takes the form of Eq. (51) and vary only the value of $E_{b}$; however, the elasticity distribution can in fact be more sophisticated and involve more terms (and our dynamics model and Julia-based simulation code can readily adapt to that by changing a single line of function definition), and thus more study can be done on the elasticity distribution part as well.
The parameter sweep on the body length implies that keep increasing the body length is not necessarily beneficial for either speed or efficiency. Although one can obtain more total thrust since a longer body can provide a larger added mass in total, the long body can also bring more burdens since it has more total mass and drag. Moreover, the optimal body stiffness distributions at different body lengths can vary significantly, which is worth investigating in further parameter studies.
Throughout our design analysis, we have been sweeping over parameters of interest respectively. Although this method can provide us with preliminary insights on fish swimming and potential design of robotic fish, the optimizations are not performed in the full parameter space. In addition, our current parameter-sweep method does not fully utilize the automatically differentiable characteristic of our dynamics model. Therefore, an important next step is to utilize the differentiability of our model and implement numerical optimization algorithms, such as gradient descent and quasi-Newton methods, to optimize propulsion performances.
VII Conclusion and Future Work
@@
To conclude, our work presents a physics-grounded and interpretable model for the full-body dynamics of fish swimming. Our model captures the dynamics of the distributed elasticity of the fish body under large deformations, while also incorporating reactive and resistive hydrodynamic forces, offering a continuous and computationally tractable alternative to multi-segment, CFD, or FSI-based approaches. A preliminary parameter study on how actuation frequency, body stiffness, and body length affect speed and propulsion frequency offers key insights into biological swimming and the design of robotic swimmers.
@@
Our future work includes utilizing the automatically differentiability of our dynamics model to perform high-dimensional numerical optimizations over the full parameter space to find the optimal geometries and material properties for maximizing swimming speed and efficiency. Moreover, we are interested in further model reductions to enable real-time model-based controls with hardware implementations on a robotic fish.

### v2

# Modeling Elastic-Body Dynamics of Robotic Fish Using a Variational Framework

@@ Abstract
Fish-inspired aquatic robots are gaining increasing attention in marine robot communities due to their high swimming speeds and efficient propulsion enabled by flexible bodies that generate undulatory motions. To support the design optimization and control of such systems, accurate, interpretable, and computationally tractable modeling of the underlying swimming dynamics is indispensable. In this letter, we present a full-body dynamics model for motor-actuated robotic fish, rigorously derived from Hamilton’s principle. The model captures the continuously distributed elasticity of a deformable fish body undergoing large deformations and incorporates fluid–structure coupling effects, enabling self-propelled motion without prescribing kinematics. Preliminary open-loop simulations examine how actuation frequency and body stiffness influence the swimming speed and energy efficiency of the robotic fish. Closed-loop simulations further assess how stiffness distribution impacts the controller’s velocity-tracking performance and energy efficiency. The results demonstrate the model’s potential for performance evaluation and control optimization of soft robotic swimmers when stiffness is treated as a design variable.
@@ I Introduction
Biomimetic swimming robots have drawn increasing interest in marine robototics, due to their potential for energy-efficient and agile locomotion in complex aquatic environments [1, 2]. Among these biomimetic swimmers, fish-like robots are particularly promising, since they utilize flexible bodies and undulatory motion to achieve high speed, propulsive efficiency, and maneuverability [3, 4, 5]. As their performance strongly depends on the coupled dynamics between body deformation and hydrodynamic forces, accurate dynamics modeling is indispensable – not only for achieving reliable simulation and control, but also for guiding design optimization and performance enhancement of bio-inspired robotic swimmers.
Plenty of theoretical and computational efforts have been made to study the dynamics of robotic fish. A common approach is to model the dynamics based on multi-segment methods [6, 7, 8, 9, 10, 11]. Most of these methods model the fish as a chain of rigid bodies, and the elasticity of the fish’s body is lumped into the joints using torsional springs; though the work in [11] adopts an elastic rod framework, the actual implementation reduces the fish body to a chain of rigid segments with locally defined bending potentials limited to the connections of segments. An exception is the model presented in [9], where an arc is used to approximate the deformation of each segment and thus adds another degree of freedom for each segment.
@@
Another approach to studying robotic fish dynamics is to perform numerical simulations based on computational fluid dynamics (CFD) and fluid-structure interactions (FSI); the latter solves for both the solid and fluid fields. The FSI-based simulations presented in [12] successfully provides high-fidelity results of muscle-actuated fish swimming. A CFD-based model that accurately captures the vortex shedding during fish swimming is presented in [13]; however, the work prescribes the deformation pattern of the fish’s body, and thus does not provide a full-dynamics analysis. The work in [14] also presents an FSI-based model that provides a detailed analysis on the fluid dynamics of surrounding water and the effects of stiffness tuning on propulsion performances; however, the solid part of the model is reduced to a two-rigid-body linkage with a joint loaded by nonlinear spring and damper.
@@
Building on the preceding discussion, this letter presents a physics-grounded and fully interpretable model that captures the full-body dynamics of fish swimming. The model accounts for continuously distributed body elasticity with arbitrarily large deformations, incorporates both reactive and resistive hydrodynamic forces from the surrounding fluid, and enables self-propulsion without any imposed or prescribed kinematics. We further conduct parameter studies based on numerical simulations. Preliminary open-loop simulations reveal how actuation frequency and body stiffness influence propulsion speed and efficiency; closed-loop simulations further investigate the relationship between stiffness distribution and the accuracy and energy efficiency of the controller.
@@ IV-C The Projected Weak Form
It is worth reiterating that all spatial integrals are evaluated using the Gauss-Legendre quadrature, and all derivatives (gradient, Hessian, and Jacobian computations) are evaluated using AD. The resulting ODEs can then be solved using numerical integrators that are accessible in most scientific computing languages.
IV-D Algorithm Implementation
We implement our numerical simulations in Julia, using the ForwardDiff.jl package for automatic differentiations and the DifferentialEquations.jl package for numerical integrations [25]. Since the Jacobian matrix of this system is naturally ill-conditioned given the strong couplings, stiffness-aware methods (e.g., implicit methods) are needed for both efficiency and accuracy. We mainly use two integrators from the package - RadauIIA5 (a stable and accurate solver that implements the implicit Runge-Kutta method) and TRBDF2 (an efficient solver that combines the trapezoidal rule with the second-order backward difference formula). Our Julia code runs on a Dell G15 laptop with a 13th Gen Intel(R) Core(TM) i7-13650HX 2.60 GHz CPU and 16 GB RAM.
@@ V Results
To validate our model, and to provide a framework for the design and feedback control of the robotic fish, we first perform a convergence experiment for the Ritz series, and then implement design analysis based on numerical simulation results. We test performances with both open-loop and closed-loop control, revealing how speed, energy efficiency, and velocity-tracking performance are influenced by actuation frequency and stiffness distribution.
In the open-loop design analysis, we sweep over two parameters respectively - the tail-beat frequency and the baseline Young’s modulus of elasticity of the fish’s body, where the Young’s modulus distribution is modeled as
@@
where $E_{b}$ is the baseline Young’s modulus. Although our model can accommodate any sufficiently well-behaved function describing the elasticity distribution along the robotic fish’s body, we use the cubic function to mimic the decrease of the body stiffness near the tail location, as this type of distribution is shown to provide efficient propulsion in robotic fish design [26]. The amplitude of the joint motor oscillations is fixed at $0.4\text{\,}\mathrm{rad}$ (i.e., the joint motor sweeps over a range of $45.86\text{\,}\mathrm{\SIUnitSymbolDegree}$). We focus on two criteria for evaluating the propulsion performances of the robotic fish - the forward speed and the cost of transport (COT), defined as
@@
where $W_{m}$ is the total work done by the joint motor, $m_{\text{total}}$ is the total mass of the robotic fish, $g$ is the gravitational acceleration, and $d$ is the total distance travelled. Since the fish undulates as it moves forward, we use the displacement magnitude in the forward direction (i.e., $X$-direction) as the effective travelled distance. The COT serves as an important metric for evaluating the efficiency of propulsion during fish swimming.
In the numerical expriments with closed-loop velocity control, we test with different baseline body elasticities and different target velocities, where the tracking performance and the COTs at different elasticities and control goals are evaluated.
During all simulations, we fix the body length to $0.3\text{\,}\mathrm{m}$ and use a PD controller on the joint angle to control the frequency and amplitude of the joint motor.
@@ V-A Convergence of Ritz Series
The correct number of Ritz basis functions (i.e., the number of terms used in the Ritz series) at which the system shows convergent behavior need to be chosen. We set the baseline Young’s modulus to $0.35\text{\,}\mathrm{MPa}$, a typical Young’s modulus for bio-based soft materials [27], and set the tail-beat frequency to $2\text{\,}\mathrm{Hz}$ and joint motor amplitude to $0.4\text{\,}\mathrm{rad}$ (i.e., the joint motor sweeps over a range of $45.86\text{\,}\mathrm{\SIUnitSymbolDegree}$). We increase the number of Ritz basis functions from 1 to 6, and compare the simulated trajectories over 3 full tail-beats (i.e., $1.5\text{\,}\mathrm{s}$) from rest and the time consumed by the numerical simulations. The comparisons are shown in Fig. 3.
@@
V-B Open-Loop Parameter Sweep on Tail-Beat Frequency
We fix the baseline Young’s modulus to $0.35\text{\,}\mathrm{MPa}$ and perform a parameter sweep over the tail-beat frequency, ranging from $0.5\text{\,}\mathrm{Hz}$ to $3.5\text{\,}\mathrm{Hz}$ with an interval of $0.5\text{\,}\mathrm{Hz}$; the results are shown in Fig. 4.
@@ V-B Open-Loop Parameter Sweep on Tail-Beat Frequency
It can be observed from Fig. 4 that the steady-state forward speed peaks at a tail-beat frequency of $3\text{\,}\mathrm{Hz}$, while the COT is minimized at a tail-beat frequency of $1\text{\,}\mathrm{Hz}$. Moreover, for a tail-beat frequency higher than $1\text{\,}\mathrm{Hz}$, increasing the steady-state speed results in an increase in COT (which implies a decrease in propulsion efficiency), and vice versa. Therefore, trade-offs need to be made and priorities between speed and efficiency need to be taken.
V-C Open-Loop Parameter Sweep on Baseline Elasticity
We fix the tail-beat frequency to $2\text{\,}\mathrm{Hz}$ and perform a parameter sweep for the baseline Young’s modulus of the fish’s body, ranging from $0.2\text{\,}\mathrm{MPa}$ to $0.5\text{\,}\mathrm{MPa}$ with an interval of $0.05\text{\,}\mathrm{MPa}$; the results are shown in Fig. 5.
@@ V-C Open-Loop Parameter Sweep on Baseline Elasticity
From Fig. 5, the forward speed is optimized at $0.25\text{\,}\mathrm{MPa}$ and COT is optimized at $0.3\text{\,}\mathrm{MPa}$. Although varying the baseline Young’s modulus at the range of $0.2\text{\,}\mathrm{MPa}$ to $0.5\text{\,}\mathrm{MPa}$ has limited impact on both the steady-state speed and the COT, this does not imply that the impact of the Young’s modulus distribution throughout the fish’s body is trivial - in our numerical experiments, we assume that the Young’s modulus distribution takes the form of Eq. (51) and vary only the value of $E_{b}$; however, the elasticity distribution can in fact be more sophisticated and involve more terms. Moreover, as shown in the next subsection, even a small change in the baseline elasticity can lead to significant differences in closed-loop control performances.
V-D Performance with Closed-Loop Control
We build a feedback controller where the robotic fish’s forward speed is controlled by the tail-beat frequency. Since the instantaneous velocity of the fish’s head oscillates with the fish’s undulations, we implement an exponential moving average (EMA) filter with dynamics
$\frac{d}{dt}v_{f}=\frac{1}{\tau_{f}}(v_{i}-v_{f})$ (53)
where $v_{f}$ is the filtered forward speed, $v_{i}$ is the instantaneous forward speed as the input to the filter, and $\tau_{f}$ is the time constant of the filter. The EMA filter is a first-order LTI system with cutoff frequency $\omega_{c}=\frac{1}{\tau_{f}}$. Given that the motor’s actuation frequency is generally above $0.5\text{\,}\mathrm{Hz}$, we set the cutoff frequency of the filter to be a half of this value as $0.25\text{\,}\mathrm{Hz}$, and thus $\tau_{f}\approx$0.64\text{\,}\mathrm{s}$\$. With the filtered velocity, we then use a PI control law for the robotic fish’s forward velocity.
We test the robotic fish’s response to two different step inputs in the reference velocity from rest. For each step input, we test baseline Young’s moduli of $200\text{\,}\mathrm{kPa}$ and $350\text{\,}\mathrm{kPa}$. The results are shown in Fig. 6
Figure 6: Simulation results of the closed-loop control test with target forward speeds of (a) $0.2\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$ and (b) $0.1\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$.
From Fig. 6, for a target velocity of $0.2\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$, the $200\text{\,}\mathrm{kPa}$ baseline elasticity offers smoother tracking performance, with less oscillations, overshoots, and steady-state error; however, for a target velocity of $0.1\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$, the $350\text{\,}\mathrm{kPa}$ baseline elasticity performs better with smaller steady-state error.
As for COT, the $200\text{\,}\mathrm{kPa}$ baseline elasticity has a COT of 0.063 at $0.1\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$ and 3.629 at $0.2\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$, while the $350\text{\,}\mathrm{kPa}$ baseline elasticity has a COT of 3.295 at $0.1\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$ and 0.124 at $0.2\text{\,}\mathrm{m}\mathrm{/}\penalty 50\mathrm{s}$. This result reveals that robotic fish navigating at varying speed levels and efficiency requirements can benefit considerably from tunable stiffness, and a promising way to perform such tuning is via optimization algorithms based on our dynamics model.
VI Conclusion and Future Work
@@
To conclude, our work presents a physics-grounded and interpretable model for the full-body dynamics of fish swimming. Our model captures the dynamics of the distributed elasticity of the fish body under large deformations, while also incorporating reactive and resistive hydrodynamic forces, offering a continuous and computationally tractable alternative to multi-segment or FSI-based approaches. A preliminary parameter study reveals how actuation frequency and body stiffness affect speed and efficiency, and how body stiffness distribution can affect the accuracy and energy efficiency of closed-loop control at different target speeds.
@@
Our future work includes utilizing the automatically differentiable nature of our dynamics model to perform high-dimensional numerical optimizations over the full parameter space to find the optimal geometries and material properties for maximizing swimming speed and efficiency. Moreover, we are interested in further model reductions to enable real-time model-based controls with hardware implementations.